OpenStudy (anonymous):

Use the quadratic formula to find the roots of the equation. x^2 + 3x - 18 =0 ... i tried using formula and couldnt get it.

OpenStudy (lgbasallote):

let me see what you did with the formula..

OpenStudy (anonymous):

how do you use draw option

OpenStudy (anonymous):

you click and draw! no eraser tho ;_;

OpenStudy (anonymous):

well i plugged all the numbers in and i guess i messed somethin up

OpenStudy (lgbasallote):

just show me your substitution

OpenStudy (anonymous):

x= -3 + (radical) 3^2 - 4 (1) (-18) ------------------------------------- 2 (1)

OpenStudy (lgbasallote):

hmm i see nothing wrong there

OpenStudy (anonymous):

but the answer didnt come out right

OpenStudy (anonymous):

its a multilpe choice also

OpenStudy (lgbasallote):

\[x = \frac{-3 \pm \sqrt{3^2 - 4(-18)}}{2}\] \[x = \frac{ -3 \pm \sqrt{9 +72}}{2}\] \[x = \frac{-3 \pm \sqrt{81}}{2}\] do you see it now?

OpenStudy (anonymous):

ohhh i c what idid wrong...

OpenStudy (anonymous):

thanks

OpenStudy (lgbasallote):

<tips hat>

OpenStudy (anonymous):

<steal hat>